The perimeter of a rectangle is 76 meters. the length is 16 meters less than twice the width. find the length of the rectangle.
Answer by praseenakos@yahoo.com(507)

QUESTION:

The perimeter of a rectangle is 76 meters. the length is 16 meters less than twice the width. find the length of the rectangle.

ANSWER:

Perimeter of a rectangle is given by the formula,

P = 2 ( length + width)

That is ,
P = 2 ( L + W ) -----------------------(1)

Here we have,

P = 76 meters.

Length is given as 16 meters less than twice the width( that is lenth is given in terms of width)
So let's take,

W = x meters.

Twice of width = 2*x = 2x meters.

So we can take length as,

L = (2x - 16) meters


Now substitute all these values in equation (1)


==> 76 = 2 ( 2x - 16 + x )


==> 76 = 2 ( 2x + x -16)

==> 76 = 2 ( 3x -16 )

Now divide both sides of the equation by 2

==> 76/2 = { 2 ( 3x -16 )}/2


==> 38 = 3x - 16

Add 16 to both sides,

==> 38 + 16 = 3x -16 + 16

==> 54 = 3x

Now again divide both sides by 3

==> 54/3 = 3x/3


==> 18 = x

Thus we got the value of x which is the width of the rectangle.

W = 18 meters

==> L = 2x - 16 = 2*18 - 16 = 36 - 16 = 20 meters.


So length of the rectangle is 20 meters.


You can check your answer by giving values of L and W in equation (1) and find whether is is equal to 76 or not.

Hope you understood.


Regards.

praseenakos@yahoo.co.in